Find y in terms of x: x^-1 + 4x^-2 + 4x^-3=x^-3y
Answer by lwsshak3(11628)

Find y in terms of x: x^-1 + 4x^-2 + 4x^-3=x^-3y
multiply each term by x^3
x^2+4x+4=y
y=(x+2)^2